Accelerator Update for February 13, 2008 - February 15, 2008
Wednesday February 13
The day shift began with Operations monitoring the Tevatron (TeV) store 5903 (60.25E30) and the Recycler stash (307.3E10), with the Antiproton Source stacking (36.8E10), with the Main Injector (MI), Linac, and Booster operating normally, with the Collider Detector at Fermilab (CDF) and D-Zero (D0) using the colliding beams, and with the NuMI and SciBooNE / MiniBooNE (BNB) experiments taking beam.
At 9:02 AM, a Booster RF station (BRF18) tripped off for the first of two times in 15 minutes. It reset.
Operations terminated store 5903 at 10:17 AM.
Operators prepared CDF and D0 for collision hall accesses at 10:28 AM.
A MI expert reported at 10:30 AM that he was inserting the MI-30 collimators for the next TeV shot.
Operations reported at 10:34 AM that they had received approval to send beam to MTest whenever they were ready for it.
MI experts reported at 11:20 AM that they were going to try to increase their intensity.
At 11:21 AM, IRF8 began tripping off. Operators contacted an expert.
CDF experimenters reported at 11:25 AM that they were turning off their solenoid for about an hour to perform a special procedure.
At 11:37 AM, BRF10 tripped off. It reset.
MI experts reported at 11:48 AM that they were going to try to increase the intensity again.
At 11:51 AM, IRF8 continued to trip off.
At 11:59 AM, Operations reported that they were disabling the SY beam for shot setup. The SY experts had not gotten the beam beyond M03 due to a beam stop problem. Experts investigated and found that the drain lines for an air dryer had frozen. This meant that there wasn’t enough air pressure to keep the beam stop open. FESS mechanics were called to warm up the drain lines.
MI experts halted their study to increase beam intensity at 12:11 PM.
CDF experimenters reported at 12:21 PM that their solenoid was back up and running.
At 12:31 PM, a Pbar expert reported that a tunnel access would be needed to fix DRF1-3.
At 12:54 PM, KRF3 tripped off. It reset.
D0 and CDF were both out of access by 12:56 PM.
Operations put the TeV through a dry squeeze at 1:17 PM.
Operations put the TeV into shot setup at 1:57 PM.
At 2:04 PM, KRF7 tripped off. It reset.
At 2:16 PM, Operations reported problems with Booster beam. Operators contacted experts.
Operations reported at 4 PM that the Recycler had stashed antiprotons three times during the day shift.
Operations established store 5905 at 4:21 PM with an initial luminosity of 225.74E30.
SY experts began to tune beam to the MTest experiment at 4:36 PM.
Operations disabled the MTest beam at 6:03 PM.
At 10:49 PM, LRF2 tripped off. It reset.
At 11:39 PM, Operations reported that the TeV luminosity dropped. Operators investigated and contacted experts.
Operations reported at midnight that the Recycler had stashed four times during the evening shift.
Thursday February 14
The midnight shift began with Operations monitoring store 5905 (74.35E30) and the stash (307.3E10), with Pbar stacking (36.8E10), and with NuMI and BNB taking beam.
At 1:14 AM, BRF9 tripped off. It reset.
A TeV expert reported at 1:33 AM that he thought the loss of luminosity was due to the orbit stabilization program; it seemed to have moved the vertical correctors by an amp.
At 2:19 AM, BRF10 tripped off for the first of four times in about two hours. It reset.
At 3:01 AM, KRF5 tripped off. It reset.
At 3:03 AM, KRF3 tripped off. It reset.
Operations established beam to MTest at 6:10 AM.
Operations reported at 7:53 AM that the Recycler had stashed six times during the midnight shift.
At 8:13 AM, BRF10 tripped off. A Booster expert bypassed the station and brought BRF19 on in its place.
An RF expert brought BRF10 back on line at 8:48 AM.
At 8:41 AM, BRF9 tripped off. It reset.
TeV experts began an orbit smooth at 9:44 AM.
At 10:17 AM, BRF10 tripped off. It reset.
Operations terminated store 5905 at 10:19 AM.
At 10:38 AM, BRF3 tripped off on a ground fault indication. It reset, but wasn’t operating normally. Operators contacted experts.
Operations reported at 10:40 AM that operators had replaced a control card for a TeV corrector magnet power supply (VF43).
At 11:27 AM, BRF10 tripped off. It reset.
At 11:42 AM, BRF1 & 10 tripped off. They reset.
The Neutron Therapy Facility reported at 12:16 PM that they had completed treatment on their last patient for the day.
At 12:49 PM, BRF10 tripped off and would not reset. Operators brought BRF19 on in its place.
At 12:51 PM, Operations reported that the Linac 55 LCW system return pressure dropped. An operator contacted an expert.
At 1:42 PM, BRF9 tripped off. It reset.
A MI expert moved the MI-30 collimators into position for a TeV shot at 2:04 PM.
At 2:51 PM, BRF18 tripped off. It reset, but then developed modulator issues. Operations lowered the MI intensity so experts could work on the station.
At 3:14 PM, the Pelletron tripped off due to bad beam positions in MI. The MI beam permit tripped off at the same time. Everything reset.
Operations reported at 4:04 PM that the Recycler had stashed two times during the day shift.
At 4:32 PM, a TeV power supply (T:O2) tripped off on an “Upstream Quench” bit. There was no quench. Experts reported that they weren’t sure what was causing the problem. They continued their investigation.
At 4:40 PM, BRF9 tripped off. It reset.
At 5:48 PM, a Pbar Debuncher vacuum pump (D:IP403) tripped off. It reset.
At 5:49 PM, LRF5 tripped off. It reset.
The MTest experimenters reported at 6:11 PM that they were done with the beam for the night.
Operations put the TeV into shot setup at 6:12 PM.
At 7:24 PM, BRF10 tripped off for the first of two times in a little over an hour. It reset.
Operations established store 5909 at 8:26 PM with an initial luminosity of 221.78E30.
At 10:36 PM, the MI “QD” (quadrupole defocusing) bus tripped off on a ground fault. At the same time in the TeV, a power supply (C:S5A3A) tripped off and switched its polarity. Operators contacted experts.
At 11:13 PM, a MI expert recommended operators raise the ground fault limit and try to run with the soft fault.
At 11:18 PM, a TeV expert instructed operators on how to get C:S5A3A back on, but it wouldn’t work. This supply will be left off until the end of the store.
At 11:40 PM, BRF13 tripped off for the first of two times in three minutes. It reset.
Operations reported at 11:40 PM that they were turning MI on and taking Pbar out of standby.
Operations reported at 11:55 PM that MIRF12 had been tripping off. Operators bypassed the station.
Operations reported at midnight that the Recycler had stashed once during the evening shift.
Friday February 15
The midnight shift began with Operations monitoring store 5909 (100.65E30), the stash (133.6E10), and the stack (45.0E10), with NuMI off due to the MI “QD” ground fault problem, and with BNB taking beam.
At 12:08 AM, the MI power supply interlock loop tripped off on a “QD” ground fault.
At 12:10 AM, LRF2 tripped off for the first of many times over four minutes.
Operations reestablished the MI ramp at 12:29 AM.
Pbar resumed stacking at 12:46 AM.
At 12:49 AM, LRF2 tripped off. It reset.
An operator brought MIRF12 back on line at 1:09 AM.
At 1:12 AM, KRF7 tripped off. It reset.
At 1:17 AM, MIRF12 tripped off and continued to trip off. Operators bypassed the station.
Operations established beam to MTest at 6:05 AM.
A Pbar expert reported at 6:31 AM that he was doing some target checks and taking some orbit data.
At 7:13 AM, a vacuum pump for LRF3 tripped off. It reset.
At 7:48 AM, a Booster vacuum pump (IPRF23) tripped off and brought down BRF13 & 14. Everything reset.
Operations reported at 7:59 AM that the Recycler had stashed five times during the midnight shift.
The Plans for Friday and the Weekend
The plans for today and the weekend are to stack, stash, and store.
Comments and Suggestions
|last modified 02/18/2008 email Fermilab|
|Security, Privacy, Legal|